OpenStudy (anonymous):

can u help me solve this equation im 13 the equation is 3y+17=-13

OpenStudy (lstrasmann):

You have to get y separated from 17. So you subtract: \[3y=-13-17\]which is \[3y=-30\]now divide\[\frac{ 3y }{ 3 } = \frac{ -30 }{ 3 }\] and get \[y = -10\]